# Multi-Agent Path Finding under Limited Communication Range Constraint via Dynamic Leading

### v1

@@ Abstract
This paper proposes a novel framework to handle a multi-agent path finding problem under a limited communication range constraint, where all agents must have a connected communication channel to the rest of the team. Many existing approaches to multi-agent path finding (e.g., leader-follower platooning) overcome computational challenges of planning in this domain by planning one agent at a time in a fixed order. However, fixed leader-follower approaches can become stuck during planning, limiting their practical utility in dense-clutter environments. To overcome this limitation, we develop dynamic leading multi-agent path finding, which allows for dynamic reselection of the leading agent during path planning whenever progress cannot be made. The experiments show the efficiency of our framework, which can handle up to 25 agents with more than 90% success-rate across five environment types where baselines routinely fail.
Introduction
We want a team of agents navigate through an obstacle-rich environment to goals while maintaining constant team communication: a spanning tree created from range-limited communication between pairs of agents. This problem is relevant to scenarios like supply delivery during disasters or monitoring hostile environments, where the risk of losing agents is significant. To mitigate this risk, agents must ensure that the team is in constant communication throughout their movement. Maintaining the spanning tree while the agents head in different directions with varied lengths of actions makes pathfinding challenging in continuous time and space, even for holonomic agents. The challenge is compounded by agents starting at random positions, needing to pass through narrow passages and non-convex spaces without collisions, and then reach random goals.
The problem can theoretically be solved using a centralized approach [1, 2, 3, 4, 5], in which planning selects between team actions, each simultaneously specifying an action for all agents at once. However, this approach quickly suffers from the curse of dimensionality, as the difficulty of planning increases exponentially with the number of agents, making planning intractable for even relatively small problems. The platooning leader-follower approach [6, 7, 8, 9] mitigates this challenge by planning each agent in sequence, where one agent is selected as the leader and others act as followers, who plan so as to maintain communication to the agent that planned before them. However, establishing a fixed planning order for leader agent can result in issues during planning. These issues include the follower agents finding no action to follow the leader within communication range or becoming stuck once the leader reaches its goal (Fig. 1.a), or the team needing to spread out for the followers to reach their goals while maintaining communication range (Fig. 1.b).
@@ Introduction
To address these challenges, this work develops a framework for solving multi-agent pathfinding problem with limited communication range (MALCR), ensuring a connected communication channel in the team at any time. We propose a novel technique: dynamic leading which allows any agent to become the new leader during multi-agent tree expansion, allowing planning to further expand the multi-agent tree if the current leader cannot make progress. Once the leader agent has planned, follower agents plan, ensuring that the resulting path is always in communication to at least one agent.
@@
We introduce an algorithm for planning in multi-agent systems with communication distance constraints: MA-DL. The experiments show that our framework results in fast and effective planning, handling up to 25 agents with more than 90% success-rate across five environment types where baselines, centralized planning with composite states and platooning leader-follower approach, routinely fail.
Figure 2: An illustration of one step expanding the MATree with 3 agents. $a_{1}$, $a_{2}$, and $a_{3}$, start at $s_{1}$, $s_{2}$, and $s_{3}$, aiming for their goals $g_{1}$, $g_{2}$, and $g_{3}$ while avoiding the obstacles (black) and other agents. From node $v_{1}$ in the multi-agent tree $\mathcal{T}_{ma}$ (step 1), the planner implements single plannings with random planning order (step 2–5), then expand $\mathcal{T}_{ma}$ based on the paths found (step 6). In step 3, $a_{2}$ follows $a_{3}$ to maintain communication links. At $t_{1}$(F), when $a_{3}$ stops, $a_{2}$ switches to follow $a_{1}$. Our key contribution, dynamic leading, occurs at step 4 when $a_{3}$ follows $a_{2}$ to $t_{2}$ (L), then becomes the new leader and plan to $t_{3}$. In the next iteration, other agents follows $a_{3}$ before expanding the MATree in step 6.
Related Work
@@
There exist multiple approaches that seek to solve multi-agent path finding problems, in the absence of a communication constraint, that could in theory be adapted to solve the MALCR problem. To effectively scale to more agents, approaches in this space are typically designed around reducing the exponential growth of the planning tree with planning time horizon. Wagner and Choset [1] proposed M algorithm, which plans individually for each agent, yet builds composite states for agents and expands the local planning tree when their paths are in collision. Though M works well if the agents are far apart, when agents are close to each other, collisions occurs frequently, causing the planning joint-state increases rapidly until fully incooperating of all agents and limiting its efficacy for solving our limited-communication problem, which requires that agents be near to one another. Solovey, Salzman, and Halperin [2] introduced the dRRT planner, which uses a implicit composite roadmap—a product of single-agent roadmaps. Their algorithm grows a multi-agent tree in this composite roadmap using RRT. However, using such a roadmap requires abstracting low-level details of each robot’s motion over time, making it difficult to apply their approach for planning with communication constraints. PIBT [5] and BMAA* [3] expand a single action for each agent in each planning iteration. These planners can handle well deadlocks among the agents. However, there are two issues for these planners to solve MALCR problem. The first one is their short-term action selection strategy, which can result in situations where followers are unable to find actions that keep them within the leader’s communication range. The second one is their fixed planning order causing followers fail to plan when their initial positions are out of the leader’s communication range (Fig. 1b). Okumuar et al. [4] proposed the OTIMAPP planner which employs Deadlock-based Search (DBS) to resolve conflict among paths determined through prioritized planning. In addition to the limitation of a fixed planning order (as mentioned above), OTIMAPP requires a new technique to modify the computed paths while adhering to deadlock and communication constraints. In general, it is very challenging to adapt these approaches to solve the MALCR problem in which each member of the team must be in constant communication with one another in continuous time.
@@
Specific to the MALCR problem, platooning has emerged as a state-of-the-art solution for multi-agent path finding in which the motion of a leader-agent is planned first and follower-agents planned one at a time in sequence to follow the agent that proceeded them, and so is designed to maintain communication constraints between pairs of agents. Most existing work in platooning performs full motion planning only for the leader, and uses a low-level controller to regulate followers to maintain a distance from the preceding agents [7, 6]. Qian et al. [8], Zhao et al. [10], and Gao et al. [9] introduced model predictive controller that generate trajectories for a virtual center, which followers need to track.
@@
However, recent work in the space of conflict-based search (CBS) [11] has proven that planning with a fixed vehicle order in general, including platooning, is incomplete and frequently becomes stuck when the leader moves in a direction different to the follower’s goals, as we show in Fig. 1, or when the team must shuffle their planning order during travel to reach their goals without violating communication constraints. Overcoming this limitation in general requires a planner that allows for dynamic leading, in which the leader can change during multi-agent tree expansion when the team fails to make progress towards the goal.
Problem Definition
@@
We formally define the Multi-Agent Path Finding with Limited Communication Range (MALCR) problem as follows. There are $n$ agents $\{a_{1},\ldots,a_{n}\}$ and a known world $\mathcal{W}$ with obstacles $\mathcal{O}=\{\mathcal{O}_{1},\ldots,\mathcal{O}_{m}\}$. The agents start at the initial positions $s^{{\mathrm{init}}}={\langle{s^{{\mathrm{init}}}_{1},\ldots,s^{{\mathrm{init}}}_{n}}\rangle}$ and head to the goals $g=\{g_{1},\ldots,g_{n}\}$, where $s_{i},g_{i}\in\mathcal{W}$ and $s_{i},g_{i}\notin\mathcal{O}$. The initial positions and goals are chosen so as to satisfy the team communication constraint, formally defined below.
@@
A collision occurs between two agents, $a_{i}$ and $a_{j}$, when their distance is less than a threshold $d_{c}$ at timestep $t$. A path is a sequence of waypoints to transition an agent from a position $p_{s}$ to position $p_{g}$ with constant velocity $\emph{v}_{c}$. We say a path is reach-goal if it can lead the agent to the goal, is collision-free, and the movement between two sequential waypoints satisfies ACOMM constraint; When an agent stops at its goal, we still consider collision and ACOMM constraints. A path is valid if it is reach-goal and at the goal $p_{g}$ from $t_{g}$, the agent’s action Stop still satisfies ACOMM and is collision-free.
The objective is to compute valid paths $\{\zeta_{1},\ldots,\zeta_{n}\}$, one for each agent, so that $\zeta_{i}$ starts at $s^{{\mathrm{init}}}_{i}$, reaches $g_{i}$, and the agents satisfy TCOMM constraints from the start time to the time the last agent reaches its goal. We develop a MAPF framework that seeks to reduce the overall planning time and the travel distances.
@@
7:        ${{\mathrm{allRG}}}\leftarrow{{\texttt{true}}}$; if $k>m_{t}$ then ${\scriptstyle{{\textsc{Shuffle}}}}(p)$
@@
10:           $\zeta^{t}\!\leftarrow\!{\scriptstyle{{\textsc{SAPF}}}}(p_{i},g_{p_{i}},\zeta^{s},\mathcal{G},\mathcal{H}_{i})$; $\zeta^{s}_{p_{i}}.{\scriptstyle{{\textsc{Insert}}}}(\zeta^{t})$
@@
Methods
@@
1:   $\mathcal{G}^{s}={\scriptstyle{{\textsc{AddNodes}}}}(\mathcal{G},g_{\rho},\zeta^{s}_{\rho}.{{\mathrm{end}}})$
2:  $v_{g}\!\leftarrow\!{\scriptstyle{{\textsc{GetNode}}}}(g_{\rho},\mathcal{G}^{s})$; $v_{s}\!\leftarrow\!{\scriptstyle{{\textsc{GetNode}}}}(\zeta^{s}_{\rho}.{{\mathrm{end}}},\mathcal{G}^{s})$
3:  ${{\mathrm{openList}}}.{\scriptstyle{{\textsc{Add}}}}(v_{s})$; ${{\mathrm{closedSet}}}\leftarrow\{\}$; $v_{{\mathrm{best}}}\leftarrow v_{s}$
4:  while ${{\mathrm{openList}}}$ not empty and ${\scriptstyle{{\textsc{time}}}}()<{{\mathrm{t}}}_{{\mathrm{sa}}}$ do
5:     $v\leftarrow{{\mathrm{openList}}}.{\scriptstyle{{\textsc{Pop}}}}()$; ${{\mathrm{closedSet}}}.{\scriptstyle{{\textsc{Add}}}}(v)$
6:     if $v=v_{g}$ then return ${\scriptstyle{{\textsc{getpath}}}}(v_{s},v,\mathcal{G}^{s})$
7:     for $u$ in ${\scriptstyle{{\textsc{GetNeighbors}}}}(v)$ do
8:        if $u\textbf{ is not in }{{\mathrm{closedSet}}}$ then
9:           $u.{{\mathrm{parent}}}\leftarrow v$; $d_{uv}\leftarrow{\scriptstyle{{\textsc{Distance}}}}(v,u)$; $u.g\leftarrow v.g$ + $d_{uv}$
10:           $u.t\leftarrow v.t$ + $d_{uv}/\text{v}_{c}$; $u.h\leftarrow{\scriptstyle{{\textsc{GetHeuristic}}}}(u,\mathcal{H}_{\rho})$; $u.f=u.g+u.h$
@@
12:              ${{\mathrm{openList}}}.{\scriptstyle{{\textsc{Add}}}}(u)$
13:              if $u=v_{g}$ then return ${\scriptstyle{{\textsc{GetPath}}}}(v_{s},u,\mathcal{G}^{s})$
14:        else if $v.g+d_{uv}<u.g$ then
15:           $u_{1}\leftarrow{\scriptstyle{{\textsc{Copy}}}}(u)$; $u_{1}.t\leftarrow v.t+d_{uv}/\text{v}_{c}$
16:           if IsNodeValid($\rho,\zeta^{s},u_{1},v,r_{c}$then
17:              $u.{{\mathrm{parent}}}\leftarrow v$; $u.t\leftarrow v.t+d_{uv}/\text{v}_{c}$; $u.g\leftarrow v.g+d_{uv}$; $u.f=u.g+u.h$
18:        if $v_{{\mathrm{best}}}.f>u.f$ then $v_{{\mathrm{best}}}\leftarrow u$
19:  return ${\scriptstyle{{\textsc{GetPath}}}}(v_{s},v_{{\mathrm{best}}},\mathcal{G}^{s})$
Our framework, named as MA-DL, consists of two modules: Multi-agent Path Finding with Dynamic Leading (MA-DL) and Single-Agent Path Finding (SAPF).
High Level Overview: The main loop of the MA-DL module (Alg. 1) expands a multi-agent (MA) planning tree $\mathcal{T}_{ma}$ whose growth is illustrated in Fig. 2. The MA tree is used to represent motion expansion of all agents. Its role is to keep track of planning progress for all agents, and the node with lowest value of cost and heuristic will be selected to expand in the next iteration. To ensure all nodes in $\mathcal{T}_{ma}$ have a chance of being selected, their costs are penalized after each selection. The MA tree expansion occurs in a loop (Fig. 2:2–5), in which each agent’s path is expanded in order. The single-agent planner (SAPF, Alg. 2) finds reach-goal shortest paths for the agents one by one according to the planning order $p$, such that each agent maintains communication with at least one of the agents that plans before it. Difference to prioritized planning approach [12] with fixed leading, the leader agent in our planner is dynamic: through dynamic leading technique (Alg. 3) which allows any agent take the leading role if it goes further than other agents, or shuffling the planning order whenener planning cannot further expand the multi-agent tree (Alg. 1:6). After each planning loop, the multi-agent tree is expanded with new nodes built from the individual paths (Fig. 2.6). Our dynamic leading technique (Fig. 2.4) allows the MA-DL planner to handle the challenging situations in Fig. 1.
MA-DL MODULE
This module manages the multi-agent tree (MATree) $\mathcal{T}_{ma}$, selects nodes to expand, dynamically selects planning orders, and triggers single-agent plannings.
The module gets the number of agents $n$, the goals $g$, initial start $s^{\text{init}}$, and the world $\mathcal{W}$ as inputs and returns the valid paths or, if valid paths cannot be found, those closest to the goal. The algorithm (Alg. 1) starts by initializing a multi-agent tree $\mathcal{T}_{ma}$ with the root consisting of all initial positions $s^{\text{init}}$ (Alg. 1:1). The world $\mathcal{W}$ is divided into sub-divisions $\Delta$ of a predefined area, with obstacle-overlapping sub-divisions further subdivided along their largest dimension until they clear or reach a size threshold. A graph $\mathcal{G}$ is created from the centroids of the obstacle-free sub-divisions, while the edges represent the connections between neighboring sub-divisions (sharing boundaries or corners). We compute the shortest paths $\mathcal{H}$ from all sub-divisions to the agent goals $g$ as heuristics by the function CalHeuristic() (Alg. 1:2). In the main loop (Alg. 1:3–15), a node $v$ with smallest cost (defined at Alg. 4.14) is selected from $\mathcal{T}_{ma}$. The cost of $v$ is then increased to encourage selecting other nodes. A heuristic function INITORDER() returns an initial planning order $p$. A list of $n$ paths $\zeta^{s}$ is initialized from $v$ (Alg. 1:4). If the node $v$ is visited, the planning order p is shuffled to get a different planning order (Alg. 1:5). The planning loop (Alg. 1:6–13) attempts to find valid paths for the agents. The loop starts by setting the variable ${{\mathrm{allRG}}}$ to true (Alg. 1:7). The planning order p is shuffled after $m_{t}$ iterations of failing to reach the goal (Alg. 1:7). The for loop (Alg. 1:8–12) plans for each agent in order by calling the module SAPF and then the returned paths is inserted into $\zeta^{s}_{p_{i}}$ (Alg. 1:10). If $\zeta^{s}_{p_{i}}$ is reach-goal, it is then checked for validity (valid) by the function ${\scriptstyle{{\textsc{MODIFYIFOVERLAP}}}}()$. The path is invalid if the Stop action at the goal at time $t$ blocks the future movement of an already-planned agent (according to planning order $p$), a situation called collision-at-goal by Bui et al. [13]. If the situation occurs, the function modifies the earlier planned paths and set ${{\mathrm{allRG}}}$ to false (Alg. 1:11). If the path is not reach-goal, ${{\mathrm{allRG}}}$ is set to false (Alg.1:12).
After all agent planning has completed, the function EXPANDMATREE() is triggered to expand $\mathcal{T}_{ma}$ from $v$ using the agent’s paths $\zeta^{s}$ (Alg. 1:14); then the node $v$ is also marked as visited. If all agents reach the goals (${{\mathrm{allRG}}}$ is still true), we break the while loop (Alg. 1:15), then return paths for all agents from $\mathcal{T}_{ma}$ (Alg. 1:16–17).

### v2

@@ Abstract
This paper proposes a novel framework to handle a multi-agent path finding problem under a limited communication range constraint, where all agents must have a connected communication channel to the rest of the team. Many existing approaches to multi-agent path finding (e.g., leader-follower platooning) overcome computational challenges of planning in this domain by planning one agent at a time in a fixed order. However, fixed leader-follower approaches can become stuck during planning, limiting their practical utility in dense-clutter environments. To overcome this limitation, we develop dynamic leading multi-agent path finding, which allows for dynamic reselection of the leading agent during path planning whenever progress cannot be made. Simulation experiments show the efficiency of our planner, which can handle up to 25 agents with more than 90% success-rate across five environment types where baselines routinely fail.
1 Introduction
We want a team of agents to navigate through an obstacle-rich environment to reach goals while maintaining constant team communication: a spanning tree created from range-limited communication between pairs of agents. This problem is relevant to scenarios such as supply delivery during disasters or monitoring hostile environments, where the risk of losing agents is significant. To mitigate this risk, agents must ensure that the team is in constant communication throughout their movement. Maintaining the spanning tree while the agents head in different directions with varied lengths of actions makes pathfinding challenging in continuous time and space, even for holonomic agents. The challenge is compounded by agents starting at random positions, needing to pass through narrow passages and non-convex spaces without collisions, and then reach random goals.
The problem can theoretically be solved using a centralized approach (?, ?, ?, ?, ?, ?, ?), in which planning selects between team actions, each simultaneously specifying an action for all agents at once. However, this approach quickly suffers from the curse of dimensionality, as the difficulty of planning increases exponentially with the number of agents, making planning intractable for even relatively small problems. The platooning leader-follower approach (?, ?, ?, ?) mitigates this challenge by planning each agent in sequence, where one agent is selected as the leader and others act as followers, who plan so as to maintain communication with the agent that planned before them. However, establishing a fixed planning order for the leader agent can result in issues during planning. These issues include the follower agents finding no action to follow the leader within the communication range or becoming stuck once the leader reaches its goal (Fig. 1.a), or the team needing to spread out for the followers to reach their goals while maintaining the communication range (Fig. 1.b).
@@ 1 Introduction
To address these challenges, this work develops a planner for solving multi-agent pathfinding problem with limited communication range (MALCR), ensuring a connected communication channel in the team at any time. We propose a novel technique: dynamic leading, which allows any agent to become the new leader during multi-agent tree expansion, allowing planning to further expand the multi-agent tree if the current leader cannot make progress. Once the leader agent has planned, follower agents plan, ensuring that the resulting path is always in communication with at least one agent.
@@
We introduce an algorithm for planning in multi-agent systems with communication distance constraints: MA-DL. Simulated experiments show that our planner results in fast and effective planning, handling up to 25 agents with more than 90% success-rate across five environment types where baselines—centralized planning with composite states and platooning leader-follower approach—routinely fail.
2 Related Work
@@
There exist multiple approaches that seek to solve multi-agent path finding problems, in the absence of a communication constraint, that could in theory be adapted to solve the MALCR problem. To effectively scale to more agents, approaches in this space are typically designed around reducing the exponential growth of the planning tree with planning time horizon. Wagner and Choset (?) proposed the M algorithm, which plans individually for each agent, yet builds composite states for agents and expands the local planning tree when their paths are in collision. Andreychuck et al. (?, ?) developed CCBS—a variant of Conflict-Based Search (CBS) in continuous time—which plans each agent individually, then resolves conflicts among agents by defining constraints for the conflicting ones and re-planning. Though M and CCBS work well if the agents are far apart, when agents are close to each other as in the MALCR problem, collisions occur frequently, causing the planners to struggle to handle the exponential growth of joint-state space or numerous agent’s path conflicts; thus limiting the efficacy of these approaches for solving our limited-communication problem, which requires the agents to be close to one another. Solovey, Salzman, and Halperin (?) introduced the dRRT planner, which uses an implicit composite roadmap—a product of single-agent roadmaps. Their algorithm grows a multi-agent tree in this composite roadmap using RRT. However, using such a roadmap requires abstracting low-level details of each robot’s motion over time, making it difficult to apply their approach for planning with communication constraints.
PIBT (?) and BMAA* (?) expand a single action for each agent in each planning iteration. These planners can handle well deadlocks among the agents. However, there are two issues for these planners to solve the MALCR problem. The first one is their short-term action selection strategy, which can result in situations where followers are unable to find actions that keep them within the leader’s communication range. The second one is their fixed planning order causing followers to fail to plan when their initial positions are out of the leader’s communication range (Fig. 1b). Okumura et al. (?) proposed the OTIMAPP planner, which employs Deadlock-based Search (DBS) to resolve conflict among paths determined through prioritized planning. Wang and Botea (?) introduced the MAPP planner in environments with requirement that, at each position on a path, an agent can always find an alternative action to bypass the next location while still reaching the one after. In addition to the limitation of a fixed planning order (as mentioned above), both OTIMAPP and MAPP would require additional algorithmic changes to modify the computed paths while adhering to deadlock and communication constraints. Choudhury et al. (?) proposed an online FV-MCTS-MP planner, which utilized coordinate graphs and the max-plus algorithm to address the challenge of action space growth as the number of agents increases. However, in the MALCR problem, neighbors are dynamic, and the agent team only needs to maintain a spanning tree rather than a fully connected graph. These features complicate the setup of local payoff components, limiting the planner’s efficiency in solving the MALCR problem and needlessly restrict the space of possible solutions. In general, it is very challenging and not straightforward to adapt these approaches to solve the MALCR problem in which each member of the team must be in constant communication with one another in continuous time.
Formation control (?, ?, ?, ?) is a straightforward method to address the MALCR problem. In this approach, leader agents are planned first, followed by the paths for follower agents, which maintain predefined formations while following the leaders that ensure constant communication among the agents. However, this approach encounters challenges when the relative positions of the agents at start and goals have complex arrangements. Additionally, in obstacle-rich environments with narrow passages, it becomes difficult for this approach to preserve the formation of the agent team.
@@
Specific to the MALCR problem, platooning has emerged as a state-of-the-art solution for multi-agent path finding in which the motion of a leader-agent is planned first and follower-agents planned one at a time in sequence to follow the agent that proceeded them, and so is designed to maintain communication constraints between pairs of agents. Most existing work in platooning performs full motion planning only for the leader and uses a low-level controller to regulate followers to maintain a distance from the preceding agents (?, ?, ?). Zhao et al. (?), and Gao et al. (?) introduced a model predictive controller that generates trajectories for a virtual center, which followers need to track.
@@
However, recent work in the space of conflict-based search (CBS) (?) has proven that planning with a fixed vehicle order in general, including platooning, is incomplete and frequently becomes stuck when the leader moves in a direction different to the follower’s goals, as we show in Fig. 1, or when the team must shuffle their planning order during travel to reach their goals without violating communication constraints. Overcoming this limitation in general requires a planner that allows for dynamic leading, in which the leader can change during multi-agent tree expansion when the team fails to make progress towards the goal.
3 Problem Definition
@@
We formally define the Multi-Agent Path Finding with Limited Communication Range (MALCR) problem as follows. There are $n$ agents $\{a_{1},\ldots,a_{n}\}$ and a known world $\mathcal{W}$ with obstacles $\mathcal{O}=\{\mathcal{O}_{1},\ldots,\mathcal{O}_{m}\}$. The agents start at the initial positions $s^{{\mathrm{init}}}={\langle{s^{{\mathrm{init}}}_{1},\ldots,s^{{\mathrm{init}}}_{n}}\rangle}$ and head to the goals $g=\{g_{1},\ldots,g_{n}\}$, where $s_{i},g_{i}\in\mathcal{W}$ and $s_{i},g_{i}\notin\mathcal{O}$. The initial positions and goals are chosen to satisfy a team communication constraint, defined below.
@@
A collision occurs between two agents, $a_{i}$ and $a_{j}$, when their distance is less than a threshold $d_{c}$ at timestep $t$. A path is a sequence of waypoints to transition an agent from a position $p_{s}$ to position $p_{g}$ with constant velocity $\emph{v}_{c}$. We say a path is reach-goal if it can lead the agent to the goal, is collision-free, and the movement between two sequential waypoints satisfies ACOMM constraint; When an agent stops at its goal, we still consider collision and ACOMM constraints. A path is valid if it is reach-goal and after reaching the goal at time $t_{g}$ forever stays at the goal, satisfies ACOMM, and is collision-free.
The objective is to compute valid paths $\{\zeta_{1},\ldots,\zeta_{n}\}$, one for each agent, so that $\zeta_{i}$ starts at $s^{{\mathrm{init}}}_{i}$, reaches $g_{i}$, and the agents satisfy TCOMM constraints from the start time to the time the last agent reaches its goal. We develop a MAPF planner that seeks to reduce the overall planning time and the travel distances.
4 Methods
Our framework, called MA-DL, consists of two modules: Multi-agent Path Finding with Dynamic Leading (MA-DL) and Single-Agent Path Finding (SAPF).
Figure 2: An illustration of one-step MATree expansion with 3 agents. Agents $a_{1}$, $a_{2}$, and $a_{3}$, start at $s_{1}$, $s_{2}$, and $s_{3}$, and target their goals $g_{1}$, $g_{2}$, and $g_{3}$ while avoiding the obstacles (black) and other agents. From node $v_{1}$ in the multi-agent tree $\mathcal{T}_{ma}$ (step 1), the planner expands trajectories for each agent, initially with a planning order, (step 2–5) then expands $\mathcal{T}_{ma}$ based on the paths (step 6). In step 3, $a_{2}$ follows $a_{3}$ to maintain communication links. At $t_{1}$(F), when $a_{3}$ stops, $a_{2}$ switches to follow $a_{1}$. Our key contribution, dynamic leading, occurs at step 4 when $a_{3}$ follows $a_{2}$ to $t_{2}$ (L), then becomes the new leader and plans to $t_{3}$. In the next iteration, other agents follow $a_{3}$ before expanding the MATree in step 6.
High Level Overview: The main loop of the MA-DL module (Alg. 1) expands a multi-agent (MA) planning tree $\mathcal{T}_{ma}$ whose growth is illustrated in Fig. 2. The MA tree is used to represent motion expansion of all agents. Its role is to keep track of planning progress for all agents, and the node with lowest value of cost and heuristic will be selected to expand in the next iteration. To ensure all nodes in $\mathcal{T}_{ma}$ have a chance of being selected, their costs are penalized after each selection. The MA tree expansion occurs in a loop (Fig. 2:2–5), in which each agent’s path is expanded in order. The single-agent planner (SAPF, Alg. 2) finds reach-goal shortest paths for the agents one by one according to the planning order $p$, such that each agent maintains communication with at least one of the agents that progresses further than it. Different from prioritized planning approaches (?) with fixed leading, our planner employs a novel dynamic leading technique (Alg. 3) and planning order shuffling (Alg. 1:6) to dynamically change the leader agent whenever planning cannot further expand the multi-agent tree. Dynamic leading enables any agent in single-agent planning to assume the leading role if it progresses further than others, effectively addressing the issues illustrated in Fig 1.a-b. After each planning loop, the multi-agent tree is expanded with new nodes built from the individual paths (Fig. 2.6).
@@ 4 Methods
7:        ${{\mathrm{allRG}}}\leftarrow{{\texttt{true}}}$
@@
10:           $\zeta^{t}\!\leftarrow\!{\scriptstyle{{\textsc{SAPF}}}}(p_{i},g_{p_{i}},\zeta^{s},\mathcal{G},\mathcal{H}_{p_{i}})$; $\zeta^{s}_{p_{i}}.{\scriptstyle{{\textsc{Insert}}}}(\zeta^{t})$
@@
4.1 Multi-Agent Path Finding with Dynamic Leading: The MA-DL Module
This module manages the multi-agent tree (MATree) $\mathcal{T}_{ma}$, selects nodes to expand, dynamically chooses planning orders, and triggers single-agent plannings.
The module gets the number of agents $n$, the goals $g$, initial start $s^{\text{init}}$, and the world $\mathcal{W}$ as inputs and returns the valid paths or, if valid paths cannot be found, those closest to the goals. The algorithm (Alg. 1) starts by initializing a multi-agent tree $\mathcal{T}_{ma}$ with the root consisting of all initial positions $s^{\text{init}}$ (Alg. 1:1). The world $\mathcal{W}$ is divided into sub-divisions $\Delta$ of a predefined area, with obstacle-overlapping sub-divisions further subdivided along their largest dimension until they are clear or reach a size threshold. A graph $\mathcal{G}$ is created from the centroids of the obstacle-free sub-divisions, while the edges represent the connections between neighboring sub-divisions (sharing boundaries or corners). We compute the shortest paths $\mathcal{H}$ from all sub-divisions to the agent goals $g$ as heuristics by the function CalHeuristic() (Alg. 1:2). In the main loop (Alg. 1:3–15), a node $v$ with smallest cost (defined at Alg. 4.14) is selected from $\mathcal{T}_{ma}$. The cost of $v$ is then increased to encourage selecting other nodes. A heuristic function INITORDER() returns an initial planning order $p$. A list of $n$ paths $\zeta^{s}$ is initialized from $v$ (Alg. 1:4). If the node $v$ is visited, the planning order p is shuffled to get a different planning order (Alg. 1:5). The planning loop (Alg. 1:6–13) attempts to find valid paths for the agents. The loop starts by setting the variable ${{\mathrm{allRG}}}$ to true (Alg. 1:7). The for loop (Alg. 1:8–12) plans for each agent in order by calling the module SAPF and then the returned paths is inserted into $\zeta^{s}_{p_{i}}$ (Alg. 1:10). If $\zeta^{s}_{p_{i}}$ is reach-goal, it is then checked for validity (valid) by the function ${\scriptstyle{{\textsc{MODIFYIFOVERLAP}}}}()$. The path is invalid if the Stop action at the goal at time $t$ blocks the future movement of an already-planned agent (according to planning order $p$), a situation called collision-at-goal by Bui et al. (?). If the situation occurs, the function modifies the earlier planned paths and set ${{\mathrm{allRG}}}$ to false (Alg. 1:11). If the path is not reach-goal, ${{\mathrm{allRG}}}$ is set to false (Alg.1:12).
After all agent planning has completed, the function EXPANDMATREE() is triggered to expand $\mathcal{T}_{ma}$ from $v$ using the agent’s paths $\zeta^{s}$ (Alg. 1:14); then the node $v$ is also marked as visited. If all agents reach the goals (${{\mathrm{allRG}}}$ is still true), the while loop is broken(Alg. 1:15), then the planner returns paths for all agents from $\mathcal{T}_{ma}$ (Alg. 1:16–17).
@@ 4.1 Multi-Agent Path Finding with Dynamic Leading: The MA-DL Module
1:  $\mathcal{G}^{s}={\scriptstyle{{\textsc{AddNodes}}}}(\mathcal{G},g_{\rho},\zeta^{s}_{\rho}.{{\mathrm{end}}})$; $v_{g}\!\leftarrow\!{\scriptstyle{{\textsc{GetNode}}}}(g_{\rho},\mathcal{G}^{s})$; $v_{s}\!\leftarrow\!{\scriptstyle{{\textsc{GetNode}}}}(\zeta^{s}_{\rho}.{{\mathrm{end}}},\mathcal{G}^{s})$
2:  ${{\mathrm{openList}}}.{\scriptstyle{{\textsc{Add}}}}(v_{s})$; ${{\mathrm{closedSet}}}\leftarrow\{\}$; $v_{{\mathrm{best}}}\leftarrow v_{s}$
3:  while ${{\mathrm{openList}}}$ not empty and ${\scriptstyle{{\textsc{time}}}}()<{{\mathrm{t}}}_{{\mathrm{sa}}}$ do
4:     $v\leftarrow{{\mathrm{openList}}}.{\scriptstyle{{\textsc{Pop}}}}()$; ${{\mathrm{closedSet}}}.{\scriptstyle{{\textsc{Add}}}}(v)$
5:     if $v=v_{g}$ then return ${\scriptstyle{{\textsc{getpath}}}}(v_{s},v,\mathcal{G}^{s})$
6:     for $u$ in ${\scriptstyle{{\textsc{GetNeighbors}}}}(v)$ do
7:        if $u\textbf{ is not in }{{\mathrm{closedSet}}}$ then
8:           $u.{{\mathrm{parent}}}\leftarrow v$; $d_{uv}\leftarrow{\scriptstyle{{\textsc{Distance}}}}(v,u)$; $u.g\leftarrow v.g$ + $d_{uv}$
9:           $u.t\leftarrow v.t$ + $d_{uv}/\text{v}_{c}$; $u.h\leftarrow{\scriptstyle{{\textsc{GetHeuristic}}}}(u,\mathcal{H}_{\rho})$; $u.f=u.g+u.h$
@@
11:              ${{\mathrm{openList}}}.{\scriptstyle{{\textsc{Add}}}}(u)$
12:              if $u=v_{g}$ then return ${\scriptstyle{{\textsc{GetPath}}}}(v_{s},u,\mathcal{G}^{s})$
13:        else if $v.g+d_{uv}<u.g$ then
14:           $u_{1}\leftarrow{\scriptstyle{{\textsc{Copy}}}}(u)$; $u_{1}.t\leftarrow v.t+d_{uv}/\text{v}_{c}$
15:           if IsNodeValid($\rho,\zeta^{s},u_{1},v,r_{c}$then
16:              $u.{{\mathrm{parent}}}\leftarrow v$; $u.t\leftarrow v.t+d_{uv}/\text{v}_{c}$; $u.g\leftarrow v.g+d_{uv}$; $u.f=u.g+u.h$
17:        if $v_{{\mathrm{best}}}.f>u.f$ then $v_{{\mathrm{best}}}\leftarrow u$
18:  return ${\scriptstyle{{\textsc{GetPath}}}}(v_{s},v_{{\mathrm{best}}},\mathcal{G}^{s})$
4.2 Single-Agent Path Finding: The SAPF Module
Single-Agent Path Finding (SAPF) finds a shortest reach-goal path for an agent. This planner searches actions on $\mathcal{G}$ that satisfy the ACOMM constraint and are collision-free. With enough runtime, the planner can explore all possible paths and so is probabilistically complete. If it fails, it returns the closest-to-goal path. Agents are assumed to have the same velocity ${{\mathrm{v}}}_{c}$. The inputs are the agent’s ID $\rho$, goal position $g_{\rho}$, a graph of the world sub-divisions $\mathcal{G}$, planned paths of agents $\zeta^{s}$, and the shortest path to the goal heuristic $\mathcal{H}_{\rho}$.
To initialize, a new graph $\mathcal{G}^{s}$ is created by adding to $\mathcal{G}$: (i) the current agent position $\zeta^{s}_{\rho}.{{\mathrm{end}}}$ and (ii) the goal $g_{\rho}$, each of which connects to the centroid of the sub-division to which they belong (Alg. 2:1). The variables $v_{g}$ and $v_{s}$ are the goal and current nodes (Alg. 2:1). $v_{s}$ is then added into ${{\mathrm{openList}}}$—a priority queue data structure, which sorts its elements by their cost-to-goal. closedSet and the closest-to-goal node $v_{\text{best}}$ are also initialized (Alg. 2:2).